the sum of three numbers is 134. the second number is 5 more than the first, and the third number is 4more than three times the second. find the numbers.
Answer by Edwin McCravy(20086)

There are three equations for the three numbers.

Let x = the first number.
Let y = the second number.
Let z = the third number.

The first equation comes from this sentence:

the sum of three numbers is 134.
x + y + z = 134

The second equation comes from this sentence:

the second number is 5 more than the first
y = x + 5

The third equation comes from this sentence:

the third number is 4 more than three times the second.
z = 3y + 4

Can you solve this syetem?:

x + y + z = 134
y = x + 5
z = 3y + 4

If not, post again asking how.

Answer: x = 22, y=27, z=85

Edwin